# The Power of Linear Programming in Sponsored Listings Ranking: Evidence from Field Experiments

### v2

@@ 3.3 Computational results
In all of our experiments, we simulate each instance (a given pair of $m,n,\lambda$) 1000 times and present the average results. For Gurobi, we used the concurrent method Gurobi 10.0, which runs primal simplex, dual simplex, and barrier methods simultaneously. We set the dual feasibility tolerance to be $10^{-4}$ (for Gurobi) and the $\epsilon$ to be $10^{-4}$ (for LP-based Algorithm 2). The experiments run on the Dell PowerEdge R640 computing node with $20$GB of memory. It is noted that the running time for Gurobi listed below only involves the solving time and does not include the model building time. In practice, the time latency ($0.1$ seconds) is required for the sum of building and solving time. To exhibit the efficiency of the LP-based algorithm, we designed two experiments.
[…]
Table 1: Experiment results for multiple $\lambda$, $m=50$, and $n=500$
$\lambda$1 .2 .3 .4 .5 .6 .7 .8 .9 .925 .95 .975
@@
In the second experiment, we would like to explore the performance of the LP-based algorithm for various problem sizes, especially when the problem dimension increases. Thus, for the second experiment, by picking $\lambda=0.95$ (which is a reasonable value for $\lambda$ as shown in the field study below), we test with a range of $m\in\{10,20,50,100,200,500\}$ and $n\in\{50,100,200,500,1000,2000\}$ ($m\leq n$). As shown in Table 2, compared to Gurobi, the LP-based algorithm achieves a massive improvement, especially under large-scale problems. We see that the time improvement grows as the problem size increases, whereas the optimal gap reduces. When $m=500$, the LP-based algorithm is more than $100$x faster than Gurobi with the optimal gap to be $0.004\%$. Under the $0.1$ seconds latency requirement, we see that the LP-based solver can solve a problem with a size of at most $m=500,n=500$. Yet, under the same conditions, Gurobi can only solve one with a size of $m=100,n=100$. Overall, the optimal gap is usually negligible ($<1\%$) for reasonably sized instances.
@@
Table 2: Experiment results for various problem sizes
@@
4 Field experiments: design and data
To demonstrate the value and power of the LP-based algorithm, we conduct a randomized field experiment that assesses the overall performance of the LP-based model. In this section, we first introduce the settings and design of this field experiment. Then we present the results of the randomization test that exhibits the rigorousness of this design.
4.1 Experiment design
We conducted an A/B test and evaluated the performance of the proposed LP-based algorithm at Marketplace A’s Marketplace across four major countries in April and May 2022. The experiment started on April 27 and lasted for 19 days until May 16. There are four notable traffic channels that Marketplace A’s marketplace has developed for consumers to visit, which are the desktop website, the iOS App, the Android App, and the mobile website, respectively. In this field experiment, we conducted the experiment in the desktop website channel (called “dWeb”), which gathered the most traffic from consumers across all channels. The benchmark we compared with in the experiments is the score-based approach described in Section 2.2, which was the production model used at Marketplace A at the time of the experiments. It is noted that, since the score-based approach was first introduced in 2020, Marketplace A has been conducting extensive experiments on it to select its optimal hyperparameters. Accordingly, in our experiment, the current version of the score-based approach is believed to have reached its full potential.
The LP-based algorithm aims to generate more revenues for the marketplace while keeping the same level of relevance for the listed products as the score-based approach obtains. In order not to impair the overall relevance level, we decided to set $\lambda\in\{0.9,0.95\}$ in the field study after examining the pre-experiment data. More specifically, we set up the following three methods, including two treatment groups and one control group:
@@ 4.1 Experiment design
    LP90 (Treatment 1) For consumers assigned to this group, the displayed listings will be generated by the LP-based algorithm with hyperparameter $\lambda=0.90$.
@@
    LP95 (Treatment 2) For consumers assigned to this group, the displayed listings will be generated by the LP-based algorithm with hyperparameter $\lambda=0.95$.

### v3

@@ A.1 Computation Results
In all of our experiments, we simulate each instance (a given pair of $m,n,\lambda$) 1000 times and present the average results. For Gurobi, we used the concurrent method Gurobi 10.0, which runs primal simplex, dual simplex, and barrier methods simultaneously. We set the dual feasibility tolerance to be $10^{-4}$ (for Gurobi) and the $\epsilon$ to be $10^{-4}$ (for LP-based Algorithm 2). The experiments run on the Dell PowerEdge R640 computing node with 2020GB of memory. It is noted that the running time for Gurobi listed below only involves the solving time and does not include the model building time. In practice, the time latency ($0.1$ seconds) is required for the sum of building and solving time. To exhibit the efficiency of the LP-based algorithm, we designed two experiments.
First, due to the special structure of the formulated SLR problem, $\lambda$ plays an essential role in determining the computational cost of solving the problem. With a small $\lambda$, the relevance constraint (8) might be redundant, thus resulting in the LP-based algorithm solving extremely fast. Therefore, we would like to compare the LP-based algorithm’s and Gurobi’s performances with a wide range of $\lambda\in[0,1]$. In the first experiment, by fixing the problem size ($m=50$ slots and $n=500$ products, which is typically the largest problem size we might encounter in Marketplace A’s SLR problem) and independently uniformly generating random variables ($h,r,v\sim U(0,1)$), we set up the test for multiple $\lambda$ ranging from 0 to $1$, i.e., $\lambda\in\{.1,.2,.3,.4,.5,.6,.7,.8,.9,.925,.95,.975\}$. The optimal gap is calculated as (Optimal value of MIP $P_{0}$ - Approximated objective value of LP-based Algorithm 2)/Optimal value of MIP $P_{0}$. As indicated in Table A1, we see that excluding the cases where $\lambda$ is too small ($\lambda\leq 0.5$), the proposed LP-based algorithm achieves more than 2020x time improvement over Gurobi LP solver. In the meantime, the LP-based algorithm can achieve the optimal gap of less than $0.05\%$ for each instance.
Table A1: Experiment Results for Multiple $\lambda$, $m=50$, and $n=500$
$\lambda$ .1 .2 .3 .4 .5 .6 .7 .8 .9 .925 .95 .975
@@
In the second experiment, we would like to explore the performance of the LP-based algorithm for various problem sizes, especially when the problem dimension increases. Thus, for the second experiment, by picking $\lambda=0.95$ (which is a reasonable value for $\lambda$ as shown in the field study), we test with a range of $m\in\{10,20,50,100,200,500\}$ and $n\in\{50,100,200,500,1000,2000\}$ ($m\leq n$). As shown in Table A2, compared to Gurobi, the LP-based algorithm achieves a massive improvement, especially under large-scale problems. We see that the time improvement grows as the problem size increases, whereas the optimal gap reduces. When $m=500$, the LP-based algorithm is more than 100100x faster than Gurobi with the optimal gap to be $0.004\%$. Under the $0.1$ seconds latency requirement, we see that the LP-based solver can solve a problem with a size of at most $m=500,n=500$. Yet, under the same conditions, Gurobi can only solve one with a size of $m=100,n=100$. Overall, the optimal gap is usually negligible ($1\%$) for reasonably sized instances.
@@
Table A2: Experiment Results for Various Problem Sizes
@@
A.2 Guidance for Fine-Tuning the LP-Based Algorithm
This section provides practical guidelines for fine-tuning and implementing the LP-based algorithm. The algorithm involves only one parameter, the relevance parameter $\lambda$. As a first step, we recommend that readers estimate a suitable range of $\lambda$ for subsequent A/B testing using historical data.
For each historical impression $k$ generated by the current ranking algorithm, assume access to $\boldsymbol{v}$, $\boldsymbol{r}$, $\boldsymbol{h}$, and the corresponding ranking $\boldsymbol{X}^{k}$. These inputs allow the computation of a relevance score (see Equation 5):
$\text{Rel}(\boldsymbol{X}^{k})=\frac{\sum_{i=1}^{m}\sum_{j=1}^{n}h_{i}r_{j}\boldsymbol{X}_{ij}^{k}}{a}.$ (1)
The intuition is as follows. Substituting $\text{Rel}(\boldsymbol{X}^{k})$ into Equation 5 yields $a(\text{Rel}(\boldsymbol{X}^{k})-\lambda)\geq 0$, which simplifies to $\text{Rel}(\boldsymbol{X}^{k})\geq\lambda$. In other words, $\lambda$ specifies a lower bound on the relevance score for future impressions.
We suggest that practitioners construct the empirical distribution of $\text{Rel}(\boldsymbol{X}^{k})$ across historical impressions and select a summary statistic (e.g., the mean, median, or a chosen quantile) as a benchmark value, denoted $\lambda_{0}$. Implementing the LP-based algorithm with $\lambda_{0}$ ensures that the resulting relevance scores are, on average, comparable to those under the prior ranking algorithm. This provides a principled starting point for further tuning via A/B testing.
Building on this benchmark, readers can adjust $\lambda$ according to their strategic objectives. For those aiming to improve ranking quality relative to the current strategy, we recommend starting with $\lambda_{0}$ and testing incremental increases. Specifically, testing groups can be formed with relevance parameters $\lambda_{0},\lambda_{0}+\delta,\lambda_{0}+2\delta,\ldots$ for some user-chosen step size $\delta>0$. As shown in our earlier analysis, higher values of $\lambda$ typically yield rankings that are more relevant.
Conversely, practitioners seeking to prioritize revenue may explore decreasing $\lambda$ from $\lambda_{0}$, for example, $\lambda_{0},\lambda_{0}-\delta,\lambda_{0}-2\delta,\ldots$. Ultimately, the optimal choice of $\lambda$ should be guided by the outcomes of A/B testing, balancing the trade-off between relevance and revenue in line with organizational objectives.
A.3 Proof of Theorem 1
In order to prove Theorem 1, we first show the following lemma that gives the closed-form solution of the LP-relaxed primal problem $P_{1}$:
Lemma 1 (Closed-form solution of the LP-relaxed primal problem).
Under Assumptions˜1 and 2 the optimal solution of the LP-relaxed primal problem $P_{1}$, denoted by $\boldsymbol{X}^{LP}$, has the following closed form:
@@ A.3 Proof of Theorem 1
    Case A (Nonbinding relevance constraint). Fix $\mu=0$ with $\boldsymbol{\delta}=\boldsymbol{v}$, then if
    $\text{Rel}(\text{PrimalResponse}(\boldsymbol{\delta},m))\geq\lambda a$, then the relevance constraint is nonbinding with $\mu^{*}=0$, and the LP optimal solution is integral: $\boldsymbol{X}^{LP}=\text{PrimalResponse}(\boldsymbol{\delta},m)$.
@@
    Case B (Binding relevance constraint). The relevance constraint binds when $\mu^{*}>0$. For any dual variable $\mu$, we define its associated integral solution as
    $\boldsymbol{X}_{+}(\mu):=\lim_{t\rightarrow 0}\text{PrimalResponse}(\boldsymbol{v}+(\mu+t)\boldsymbol{r},m)$, then the optimal dual variable $\mu^{*}$ of the problem $D_{1}$ is characterized as the smallest pairwise indifference value such that its associated integral solution is feasible, i.e., :
    $\mu^{*}=\inf\{\mu\in\{0\}\cup\{\mu_{jk}\}_{(j,k)\in\mathcal{I}}:\text{Rel}(\boldsymbol{X}_{+}(\mu))\geq\lambda a\}.$ (2)
    Then for such $\mu^{*}$, let $\boldsymbol{\delta}^{*}:=\boldsymbol{v}+\mu^{*}\boldsymbol{r}$ and its sorted indices in non-increasing order as $j_{(1)},\cdots,j_{(n)}$.
    • B1 (Boundary tie between position $m$ and $m+1$). If $\delta^{*}_{j(m)}=\delta^{*}_{j(m+1)}$, then the optimal solution is fractional only in the last row:
      $\boldsymbol{X}^{\text{LP}}_{i,j(i)}=1\;\forall\;i<m,\boldsymbol{X}^{\text{LP}}_{m,j(m)}=\alpha,\boldsymbol{X}^{\text{LP}}_{m,j(m+1)}=1-\alpha,\boldsymbol{X}^{\text{LP}}_{ij}=0\text{ otherwise},$ (3)
    • B2 (Interior tie within the top $m$ positions). If $\delta^{*}_{j(s)}=\delta^{*}_{j(s+1)}$ for some $s<m$, then optimal solution is fractional only in row $s$ and $s+1$:
      $\boldsymbol{X}^{\text{LP}}_{i,j(i),}=1\;\forall\;i\in[m]\setminus\{s,s+1\},\boldsymbol{X}^{\text{LP}}_{s,j(s)}=\boldsymbol{X}^{\text{LP}}_{s+1,j(s+1)}=1-\alpha,$
      $\boldsymbol{X}^{\text{LP}}_{s,j(s+1)}=\boldsymbol{X}^{\text{LP}}_{s+1,j(s)}=\alpha,\boldsymbol{X}^{\text{LP}}_{ij}=0\text{ otherwise},$ (4)
    • B3 (No tie within the top $m$ positions). If $\delta^{*}_{j(1)}>\cdots>\delta^{*}_{j(m)}$, i.e., all top-m scores are strictly ordered, then the optimal LP solution is integral:
      $\boldsymbol{X}^{\text{LP}}_{i,j(i)}=1\;\forall\;i\in[m],\boldsymbol{X}^{\text{LP}}_{ij}=0\text{ otherwise}.$ (5)
Proof.
For Case A, it is trivial to see that when the relevance constraint of $P_{1}$ is nonbinding, then the optimal dual variable $\mu^{*}=0$. The LP optimal solution is integral and the permutation of the sorted vector $\boldsymbol{v}$ in a descending order.
For Case B, we show that the derived LP solutions in all sub-cases are (1) feasible and (2) optimal.
Feasibility Check.
  • B1 (Boundary tie between position $m$ and $m+1$).
    $\text{Rel}(\boldsymbol{X}^{\text{LP}})$ $=\sum_{i=1}^{m-1}h_{i}r_{j(i)}+\alpha h_{m}r_{j(m)}+(1-\alpha)h_{m}r_{j(m+1)}$ (6)
    $=\sum_{i=1}^{m-1}h_{i}r_{j(i)}+\lambda a-\sum_{i=1}^{m-1}h_{i}r_{j(i)}$ (7)
    $=\lambda a.$ (8)
  • B2 (Interior tie within the top $m$ positions).
    $\text{Rel}(\boldsymbol{X}^{\text{LP}})$ $=\sum_{i=1,i\notin\{s,s+1\}}^{m}h_{i}r_{j(i)}+(1-\alpha)(h_{s}r_{j(s)}+h_{s+1}r_{j(s+1)})+\alpha(h_{s+1}r_{j(s)}+h_{s}r_{j(s+1)})$ (9)
    $=\sum_{i=1,i\notin\{s,s+1\}}^{m}h_{i}r_{j(i)}+\lambda a-\sum_{i=1,i\notin\{s,s+1\}}^{m}h_{i}r_{j(i)}$ (10)
    $=\lambda a.$ (11)
  • B3 (No tie within the top $m$ positions). The proposed $\boldsymbol{X}^{\text{LP}}$ is the permutation matrix that assigns the top $m$ items to the $m$ rows in the order of $\boldsymbol{h}$. We will see below that in this subcase necessarily $\text{Rel}(\boldsymbol{X}^{\text{LP}})=\lambda a$, otherwise $\mu^{*}$ would not be dual-optimal.
Optimality Check.
We show that $\boldsymbol{X}^{\text{LP}}$ in all three subcases of Case B are optimal for $P_{1}$. We first verify the KKT conditions for the relaxed LP problem:
  1. 1.
    Dual feasibility: by definition of the dual problem.
  2. 2.
    Primal feasibility: as we have shown in the previous part of Feasibility check.